REDUCED DIMENSION ROVIBRATIONAL VARIATIONAL CALCULATIONS OF THE S_{1} STATE OF C_{2}H_{2}

\underline{P. B. CHANGALA}, J. H. BARABAN, R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139, USA; J. F. STANTON, Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, The University of Texas at Austin, Austin, Texas 78712; A. J. MERER, Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan; Department of Chemistry, University of British Columbia, B.C. V6T 1Z1, Canada.

The bending and torsional degrees of freedom in S_{1} acetylene, C_{2}H_{2}, are subject to severe vibrational resonances and rovibrational interactions, which result in the low-energy vibrational polyad structure of these modes. As the internal energy approaches that of the barrier to cis-trans isomerization, these energy level patterns undergo further large-scale reorganization that cannot be satisfactorily treated by traditional models tied to local equilibrium geometries. Experimental spectra in the region near the cis-trans transition state exhibit these complicated new patterns. In order to rationalize our near-barrier observations and predict the detailed effects of cis-trans isomerization on the rovibrational energy structure, we have performed reduced dimension rovibrational variational calculations of the S_{1} state. Our calculation uses a high accuracy ab initio potential surface and a fully symmetrized extended-CNPI group theoretical treatment of a multivalued internal coordinate system that is appropriate for bending and torsional large amplitude motions. We will discuss these results and the insights they offer on understanding both large-scale features and spectroscopic details, such as tunneling staggerings, of barrier-proximal rovibrational levels of the S_{1} state. We will also discuss spectral features by which barriers can be located and characterized in general polyatomic systems.